Laval produces lamps and home lighting fixtures. Its most popular product is a brushed aluminum desk lamp. This lamp is made from components shaped in the fabricating department and assembled in the assembly department. Information related to the 27,000 desk lamps produced annually follows.

Direct materials $ 270,000
Direct labor
Fabricating department (9,000 DLH × $22 per DLH) $ 198,000
Assembly department (16,200 DLH × $28 per DLH) $ 453,600
Machine hours
Fabricating department 14,100 MH
Assembly department 20,550 MH

  
Expected overhead cost and related data for the two production departments follow.

Fabricating Assembly
Direct labor hours 170,000 DLH 265,000 DLH
Machine hours 156,000 MH 124,000 MH
Overhead cost $ 420,000 $ 415,000

  
Required
1. Determine the plantwide overhead rate for Laval using direct labor hours as a base.
2. Determine the total manufacturing cost per unit for the aluminum desk lamp using the plantwide overhead rate.
3. Compute departmental overhead rates based on machine hours in the fabricating department and direct labor hours in the assembly department.
4. Use departmental overhead rates from requirement 3 to determine the total manufacturing cost per unit for the aluminum desk lamps.

rev: 04_05_2017_QC_CS-84897

Next Visit question map

Question15of17Total15 of 17

Prev
Answer #1

SOLUTION

1. Total direct labor hours for both departments = 170,000 + 265,000 = 435,000 DLH

Total overhead cost = $420,000 + $415,000 = $835,000

Plant wide overhead rate on the basis of direct labor hours = $835,000 / 435,000 = $1.92 per DLH

2.

Particulars Amount ($)
Total Manufacturing cost:
Direct material 270,000
Direct Labor:
Fabricating department 198,000
Assembly department 453,600
Applied overhead (25,200*$1.92) 48,384
Total manufacturing cost (a) 969,984
Total desk lamp produced (b) 27,000
Manufacturing cost per unit (a/b) $35.92

3.

Particulars Fabricating Assembly
Overhead cost of department 420,000 415,000
Allocation base Machine hours Direct labor hours
Total machine hours 156,000
Total direct labor hours 265,000
Overhead rate of department $2.69 1.57

4.

Particulars Amount ($)
Total Manufacturing cost:
Direct material 270,000
Direct Labor:
Fabricating department 198,000
Assembly department 453,600
Applied overhead :
Fabricating department (14,100*$2.69) 37,929
Assembly department (16,200*1.57) 25,434
Total manufacturing cost (a) 984,963
Total desk lamp produced (b) 27,000
Manufacturing cost per unit (a/b) $36.48